# Taming spin susceptibilities in frustrated quantum magnets: Mean-field form and approximate nature of the quantum-to-classical correspondence

### v2

@@ Abstract
In frustrated magnetism the empirically found quantum-to-classical correspondence (QCC) matches the real-space static susceptibility pattern of a quantum spin-$1/2$ model with its classical counterpart computed at a certain elevated temperature. This puzzling relation was observed via bold line diagrammatic Monte Carlo simulations in dimensions two and three. The matching was within error bars and seemed valid down to the lowest accessible temperatures $T$ about an order of magnitude smaller than the exchange coupling $J$. Here we employ resummed spin diagrammatic perturbation theory to show analytically that the QCC breaks at fourth order in $J/T$ and provide the approximate mapping between classical and quantum temperatures. Our treatment further reveals that QCC is an indication of the surprising accuracy with which static correlators can be approximated by a simple renormalized mean-field form. We illustrate this for all models discussed in the context of QCC so far, including a recent example of the $S=1$ material $\mathrm{K}_{2}\mathrm{Ni}_{2}(\mathrm{SO}_{4})_{3}$. The success of the mean-field form is traced back to partial diagrammatic cancellations.
@@
While complete answers to these questions remain elusive even in theory, impressive progress has been made on the numerical front [4, 5, 6, 7, 8, 9]. During this endeavor, in 2013, a particularly puzzling empirical observation appeared for the triangular lattice quantum $S=1/2$ Heisenberg anti-ferromagnet (AFM) [10]: The bond-resolved spin susceptibility $\chi(\mathbf{r})$ [also known as the static spin correlator, see Eq. (2)] was computed via bold line diagrammatic Monte Carlo (BDMC). For all temperatures $T\!\geq\!0.375J$ attainable, the intricate and highly featured normalized pattern $\chi(\mathbf{r})/\chi(\mathbf{0})$ (see Fig. 3 left) can be matched by correlation data obtained from the classical ($S=\infty$) vector-spin version of the same model at an empirically obtained elevated temperature $T^{(c)}\!>T$! This was dubbed the quantum-to-classical correspondence (QCC)  [^1] QCC is not to be confused with the fact that quantum phase transitions in $d$ spatial dimensions can share a universality class with a classical counterpart in $d+1$ spatial dimensions [31]..
@@
Figure 1: QCC for systems with nearest-neighbor AFM Heisenberg coupling $J$, left for triangular and pyrochlore lattice with coordination number $z=6$ and right for kagome and square lattice ($z=4$). The markers are reproduced from the BD(MC) studies of Refs. [10, 14, 15] and indicate the temperature $T^{(c)}$ at which the normalized susceptibility pattern of the classical vector-spin model matches the one of the quantum $S=1/2$ model at temperature $T$ within error bars. The solid lines show the analytical prediction for the approximate QCC obtained by numerically inverting Eq. (8), the dashed line denotes the high-$T$ limit $T^{(c)}=4T/3$.
In the years following the initial observation, QCC was found wherever BDMC was aimed at: The square- and kagome lattice in $d=2$ [10, 15], the pyrochlore lattice in $d=3$ [14] as well as $J_{1}\!-\!J_{2}$ models on the square and anisotropic triangular lattices [15]. In Fig. 1, the corresponding $T^{(c)}(T)$ for various nearest-neighbor models are reproduced from Ref. [10] (markers). Recently, QCC was also observed in a $S=1$ model for the material $\mathrm{K}_{2}\mathrm{Ni}_{2}(\mathrm{SO}_{4})_{3}$ consisting of two interconnected $d=3$ trillium lattices [16], here the quantum data was obtained from the pseudo-fermion functional renormalization group (PFFRG) [7].
As of today, the QCC remains a puzzle with overwhelming empirical evidence but no explanation. Why do the celebrated quantum fluctuations maximized for the smallest spin length $S\!=\!1/2$ merely seem to be accounted for classically by an effective heating effect? And would the QCC break down and reveal an approximate nature if a more powerful successor of BDMC could reach lower $T$ and smaller error bars? If not, Ref. [10] speculated, could the $T^{(c)}(T)$ curve be extrapolated to $T=0$ such that a classical simulation would reveal properties of highly sought-after quantum ground states?
In this letter we shed analytical light on the origin of QCC building on recent progress in spin-diagrammatics [17]. The QCC turns out to be only approximate for all finite spin lengths $S$ and dimensions $d$. We quantify QCC’s failure at order $[J/T]^{4}$, provide a closed-form expression of the (approximate) $T^{(c)}(T)$ curve, and reveal a connection between QCC and the surprising success of a renormalized mean-field (MF) ansatz for $\chi(\mathbf{r})$. We also quantitatively explain the failure of QCC for the $d=1$ Heisenberg chain at intermediate temperature, as empirically pointed out already in Ref. [10].
@@
where $\beta=1/T$ and $S_{\mathbf{r}}^{z}(\tau)$ is the Heisenberg spin operator at imaginary time $\tau$. The Larkin equation [18, 19] expresses (2) via the set of one-$J$-irreducible (static) two-legged spin-correlator diagrams $\Sigma(\mathbf{k})$,
@@
$[T\chi(\mathbf{k})]^{-1}=f+g\gamma(\mathbf{k})+\epsilon(\mathbf{k}).$ (4)
In MF approximation, $f=1/b_{1}$, $g=\beta J$ and $\epsilon(\mathbf{k})=0$. In that sense, the exact $f$ is the (dimensionless) renormalized on-site inverse susceptibility, while $g$ describes the renormalized coupling. Therefore, we will refer to the right-hand side of Eq. (4) with $\epsilon(\mathbf{k})=0$ as renormalized MF form.
Figure 3: Absolute values of normalized real space spin susceptibility for AFM $S=1/2$ Heisenberg models with $J_{1}=1$ and temperature $T$ (dots, data taken from BDMC calculation of Refs. [10, 14, 15]) compared to the renormalized MF form (circles) parameterized with temperature $T_{MF}$ (see text) optimized for best agreement. The almost perfect match for a wide range of models and for rather low $T$ shows the accuracy of the renormalized MF form in $d>1$ dimensions. For pyrochlore, the mean value of $|\chi(\mathbf{r})|$ with the same $r$ is plotted to match the presentation in Fig. 4 of Ref. [14].
@@
Approximate QCC.—To address the QCC puzzle [10] we consider classical (unit-vector) spins via the $S\rightarrow\infty$ limit. This is straightforward in our general-$S$ formalism after spin-operators in Eqns. (1) and (2) are rescaled by $1/S$. The resulting classical coupling and susceptibility are denoted with superscript (c), hence $\chi^{(c)}(\mathbf{k})^{-1}=\underset{S\rightarrow\infty}{\mathrm{lim}}S^{2}\chi(\mathbf{k})^{-1}$ with $X\!\xrightarrow{}\!X^{(c)}/S^{2}$. Note that $b_{1}$ also depends on $S$, hence the right-hand sides of Eqns. (5) to (7) simplify by the erasure of quantum effects.
@@
In Fig. 1, we plot the numerical inverse of Eq. (8) [without the unknown contribution $O(X^{4})$] for various $S=1/2$ systems (lines). Note that the lattice only enters via the coordination number $z$. Results are in good agreement with the empirical data points reproduced from the (BD)MC studies of various $z=4$ and $z=6$ models in Refs. [10, 14, 15].
@@
For all lattices considered here, the corresponding classical temperature is always higher than the quantum temperature. Intuitively, quantum fluctuations heat up the system [10]. However, this is not generally the case with the spin-dimer as a counter-example (See End Matter).
@@
Renormalized MF form of susceptibility.—The QCC is only approximate and breaks down at order $X^{4}$ where first corrections $\epsilon(\mathbf{k})$ to the renormalized MF form appear, see Eq. (7). However, this correction vanishes in the classical limit, $\lim_{S\xrightarrow{}\infty}S^{2}\epsilon(\mathbf{k})=0+{O}((X^{(c)})^{6})$. Therefore, from fourth order on, the momentum dependence of the quantum and classical $\chi(\mathbf{k})$ is inherently different, making it fundamentally impossible to extend the mapping $X(X^{(c)})$ in Eq. (8) for the full $\chi(\mathbf{k})$ to order $X^{4}$. The QCC is therefore always of approximate nature and, contrary to the speculations in Ref. [10], fails for $T\rightarrow 0$.
@@
However, it turns out that corrections $\epsilon(\mathbf{k})$ to the renormalized MF form of $[T\chi(\mathbf{k})]^{-1}$ are relatively small even for small $T/J$ (large $X$). This can be anticipated from the partial cancellation of diagrams as mentioned above, but it can also be inferred empirically by existence of the QCC even for very low $T/J$, e.g. in the pyrochlore case down to $T/J=0.1667$ [14]. Formally, the renormalized MF form of $[T\chi(\mathbf{k})]^{-1}$ is a good approximation to the exact value as long as the minimum of the first two terms $\Delta=f-gz$ (the MF gap) is large compared to $\epsilon(\mathbf{k})$, such that $\|\epsilon\|_{1}/\Delta\ll 1$. Here we use a $L_{1}$ norm in real-space, $\|\epsilon\|_{1}=\sum_{j}|\epsilon_{i,j}|$.
@@
In the following we show that for the $d=2,3$ models considered in the context of QCC the susceptibility patterns are very accurately approximated by $\epsilon(\mathbf{k})\approx 0$ well into the cooperative paramagnetic regime $T/J\gtrsim 0.1$. In Fig. 3 we match the $S=1/2$ BDMC data from various models treated in Refs. [10, 14, 15] to the renormalized MF form with the empirically optimized MF temperature $T_{MF}$ proportional to the ratio of $f$ and $g$, $T_{MF}=\frac{fb_{1}}{g}J$. For the $J_{1}\!-\!J_{2}$ models with the renormalized MF form $[T\chi(\mathbf{k})]^{-1}\simeq f+g_{1}\gamma_{1}(\mathbf{k})+g_{2}\gamma_{2}(\mathbf{k})$ (See End Matter), also $J_{2}$ is adjusted, see the labels in Fig. 3. In all instances, the renormalized MF and BDMC data fit very well, even at temperatures below which the analytical mapping based on the truncated version of Eq. (8) would yield reasonable results. This is our second main result. Importantly, this goes beyond the empirical QCC [10] since the exact susceptibilities in the classical case, to which the quantum susceptibilities were matched so far, also feature $O([X^{(c)}]^{6})$ corrections to the renormalized MF form, c.f. the discussion above.
@@
As stated by Kulagin et al. [10] the QCC does not hold in the $d=1$ Heisenberg chain. The susceptibility of this model’s classical counterpart is exactly described by the renormalized MF form with $f=\frac{3u^{2}+3}{1-u^{2}}$, $g=\frac{3u}{1-u^{2}}$ and $\epsilon(\mathbf{k})=0$ where $u=\coth(X)-1/X$ [20]: Therefore, QCC can be observed as long as $\|\epsilon\|_{1}/\Delta\ll 1$ for the quantum correlator. We calculate the susceptibility for the $S=1/2$ AFM Heisenberg chain with quantum MC [21, 22]. In Fig. 4 we show that already for $T/J=0.33$, the corrections to renormalized MF form are sizable, $\|\epsilon\|_{1}/\Delta\approx 15\%$. This means that the range of $T/J$ where QCC could be observed is significantly smaller in $d=1$ than in the $d=2,3$ models, where no violation was found at the temperatures available to BDMC. This empirical finding can be rationalized in a $\frac{1}{d}$ expansion [17] where corrections to the renormalized MF form at the MF-critical temperature scale as $\epsilon(\mathbf{k})\sim\frac{1}{d^{2}}$.
@@
Figure 4: Parameters of the inverse susceptibility in Eq. (4) for the $S=1/2$ AFM Heisenberg chain of length $L=128$. Lines are a guide to the eye. For better visibility $f,g,\epsilon$ where all multiplied by the normalization constant $\chi(r=0)=\int_{k}\chi(k)$. The breakdown of the renormalized MF form (and thus of QCC) is rooted in a non-negligible $\|\epsilon\|_{1}/\Delta$.
@@
Figure 5: Static spin structure factor $S(\mathbf{k})=\sum_{\nu,{\nu^{\prime}}}\chi_{\nu,{\nu^{\prime}}}(\mathbf{k})f_{\mathrm{Ni}^{2+}}(\mathbf{k})$ along different planes in reciprocal space, where $f_{\mathrm{Ni}^{2+}}(\mathbf{k})$ is the form factor of the $\mathrm{Ni}^{2+}$ ions. Panels (a)-(c) show classical MC results at $T=0.35J_{4}$ (top left part of each panel), PFFRG results at flow parameter $\Lambda=0.58J_{4}$ (top right), and a fit to the renormalized MF form (rMF, lower half) at $T_{\mathrm{MF}}=0.787J_{4}$. Panels (e)-(g) display line cuts along three principal directions. The classical MC and PFFRG data was taken from Ref. [16].
@@
Finally, we test the renormalized MF form for a model of a complex realistic material $\mathrm{K}_{2}\mathrm{Ni}_{2}(\mathrm{SO}_{4})_{3}$ where QCC was empirically observed by Gonzalez et al. [16]. According to ab-initio calculations [16], this $d\!=\!3$ material realizes a $S\!=\!1$ Heisenberg model on two interconnected trillium lattices with $\{J_{1},J_{2},J_{3},J_{4},J_{5}\}=\{0.066,-0.026,0.144,1,0.479\}$ with dominant $J_{4}=1$. This model was treated with the ground state PFFRG [7] with results extracted at finite Matsubara frequency cutoff parameter $\Lambda=0.67J_{4}$. The latter was argued to act (at least qualitatively) as a finite temperature $T\sim\Lambda$ [23]. Unlike BDMC, the error of PFFRG is uncontrolled. However, empirical confirmation of QCC with $T^{(c)}=0.35J_{4}$ found in Ref. [16] suggests that the susceptibility is well described by a renormalized MF form. In Fig. 5 we match the classical MC and PFFRG susceptibilities [16] along multiple planes and cuts in $\mathbf{k}$-space to a renormalized MF form at $T_{\mathrm{MF}}=0.787J_{4}$ and all $J_{m}$ unchanged. The quality could be further improved by also renormalizing the latter.
@@
Conclusion.—The approximate description of quantum spin systems with classical theories is a recurring theme [24, 25, 26, 27] and cancellations of quantum effects have been observed for various specific spin models on particular lattices [28, 29]. However, the so far empirical QCC for the spin susceptibility on general two- and three dimensional Heisenberg models [10] revisited in this letter stands out by its surprising universality and accuracy even for low temperatures $T$ down to about an order of magnitude below $J$. Using resummed spin diagrammatic perturbation theory, we showed rigorously that the QCC is only approximate and fails at order $(J/T)^{4}$. At this order, the spatial dependence of the exact susceptibility deviates from a simple renormalized MF form for $S<\infty$.
@@
The success of the renormalized MF form is surprising, since ordinary MF theory is only valid for high temperatures or lattices with large coordination numbers. For future work, it is thus an interesting question if the renormalized MF ansatz could inform the development of novel theoretical (renormalization-)methods or experimental data analysis. The latter could be especially relevant for atom-tweezer array quantum simulators where $\chi(\mathbf{r})$ should be directly accessibly due to single-site control and measurement capabilities [30]. As a first step, it would be useful to extend the analytical (and approximate) QCC mapping of Eq. (8) to higher orders in $(J/T)$ for better analytical estimation of the renormalized MF parameters.
Acknowledgment.—We thank Matías G. Gonzalez, Philip Osterholz, Johannes Reuther, Nepomuk Ritz and Manuel Weber for helpful discussions. We acknowledge support from DFG grant no. 524270816. B. Sb. acknowledges support from DFG through the Research Unit FOR 5413/1 (grant no. 465199066). B.Sch. acknowledges funding from the Munich Quantum Valley, supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.

### v3

@@ Abstract
In frustrated magnetism, the empirically found quantum-to-classical correspondence (QCC) matches the real-space static susceptibility pattern of a quantum spin-$1/2$ model with its classical counterpart computed at a certain elevated temperature. This puzzling relation was observed via bold line diagrammatic Monte Carlo simulations in dimensions two and three. The matching was within error bars and seemed valid down to the lowest accessible temperatures $T$ about an order of magnitude smaller than the exchange coupling $J$. Here, we employ resummed spin diagrammatic perturbation theory to show analytically that the QCC breaks weakly at fourth order in $J/T$ and provide the approximate mapping between classical and quantum temperatures. Our treatment further reveals that QCC is an indication of the surprising accuracy with which static correlators can be approximated by a simple renormalized mean-field form. We illustrate this for all models discussed in the context of QCC so far, including a recent example of the $S=1$ material $\mathrm{K}_{2}\mathrm{Ni}_{2}(\mathrm{SO}_{4})_{3}$. The success of the mean-field form is traced back to partial diagrammatic cancellations.
@@
While complete answers to these questions remain elusive even in theory, impressive progress has been made on the numerical front [4, 5, 6, 7, 8, 9]. During this endeavor, in 2013, a particularly puzzling empirical observation appeared for the triangular lattice quantum $S=1/2$ Heisenberg anti-ferromagnet (AFM) [10]: The bond-resolved spin susceptibility $\chi(\mathbf{r})$ [also known as the static spin correlator, see Eq. (2)] was computed via bold line diagrammatic Monte Carlo (BDMC). For all attainable $T\!\geq\!0.375J$, the intricate and highly featured normalized pattern $\chi(\mathbf{r})/\chi(\mathbf{0})$ (see Fig. 3 left) can be matched by correlation data obtained from the classical ($S=\infty$) vector-spin version of the same model at an empirically obtained elevated temperature $T^{(c)}\!>T$! This was dubbed the quantum-to-classical correspondence (QCC)  [^1] QCC is not to be confused with the fact that quantum phase transitions in $d$ spatial dimensions can share a universality class with a classical counterpart in $d+1$ spatial dimensions [31]..
@@
Figure 1: QCC for systems with nearest-neighbor AFM Heisenberg coupling $J$, left for triangular and pyrochlore lattice with coordination number $z=6$ and right for kagome and square lattice ($z=4$). The markers are reproduced from the BDMC studies of Refs. [10, 14, 15] and indicate the temperature $T^{(c)}$ at which the normalized susceptibility pattern of the classical vector-spin model matches the one of the quantum $S=1/2$ model at temperature $T$ within error bars. The solid lines show the analytical prediction for the approximate QCC obtained by numerically inverting Eq. (8), the dashed line denotes the high-$T$ limit $T^{(c)}=4T/3$.
In the years following its initial observation on the triangular lattice, QCC was found wherever BDMC was aimed at: The square- and kagome lattice in $d=2$ [10, 15], the pyrochlore lattice in $d=3$ [14] as well as $J_{1}\!-\!J_{2}$ models on the square and anisotropic triangular lattices in $d=2$ [15]. The corresponding $T^{(c)}(T)$ for various nearest-neighbor models is compiled in Fig. 1 (markers). Recently, QCC was also observed in a $S=1$ model for the material $\mathrm{K}_{2}\mathrm{Ni}_{2}(\mathrm{SO}_{4})_{3}$ consisting of two interconnected $d=3$ trillium lattices [16], here the quantum data was obtained from the pseudo-fermion functional renormalization group (PFFRG) [7].
As of today, the QCC remains puzzling with overwhelming empirical evidence but no explanation. Why do the celebrated quantum fluctuations maximized for the smallest spin $S\!=\!1/2$ merely seem to be accounted for classically by an effective heating? And would the QCC break down and reveal an approximate nature if a more powerful successor of BDMC could reach lower $T$ and smaller error bars? If not, Ref. [10] speculated, could the $T^{(c)}(T)$ curve be extrapolated to $T=0$ such that a classical simulation would reveal properties of highly sought-after quantum ground states?
In this letter we shed analytical light on the origin of QCC building on recent progress in spin-$S$-diagrammatics [17]. The QCC turns out to be only approximate for all finite $S$ and dimensions $d$. We quantify QCC’s failure at order $[J/T]^{4}$, provide a closed-form expression of the (approximate) $T^{(c)}(T)$ curve, and reveal a connection between QCC and the surprising success of a renormalized mean-field (MF) ansatz for $\chi(\mathbf{r})$. We also quantitatively explain the failure of QCC for the $d=1$ Heisenberg chain at intermediate temperature, as empirically pointed out already in Ref. [10].
@@
where $\beta=1/T$ and $S_{\mathbf{r}}^{z}(\tau)$ is the Heisenberg spin operator at imaginary time $\tau$. The Larkin equation [18, 19] expresses Eq. (2) via the set of one-$J$-irreducible (static) two-legged spin-correlator diagrams $\Sigma(\mathbf{k})$,
@@
$\beta[\chi(\mathbf{k})]^{-1}=f+g\gamma(\mathbf{k})+\epsilon(\mathbf{k}).$ (4)
In MF approximation, $f=1/b_{1}$, $g=\beta J$ and $\epsilon(\mathbf{k})=0$. In that sense the exact $f$ is the (dimensionless) renormalized on-site inverse susceptibility, while $g$ describes the renormalized coupling. We refer to the right-hand side of Eq. (4) with $\epsilon(\mathbf{k})\!=\!0$ as renormalized MF form.
Figure 3: Absolute values of normalized real-space spin susceptibility for AFM $S=1/2$ Heisenberg models with $J_{1}=1$ and temperature $T$ (dots, data taken from BDMC calculations of Refs. [10, 14, 15]) compared to the renormalized MF form (circles) parameterized with temperature $T_{MF}$ (see text) optimized for best agreement. The almost perfect match for a wide range of models and for rather low $T$ shows the accuracy of the renormalized MF form in $d>1$ dimensions. For pyrochlore, the mean value of $|\chi(\mathbf{r})|$ with the same $r$ is plotted to match the presentation in Fig. 4 of Ref. [14].
@@
Approximate QCC.—To address the QCC puzzle [10] we consider classical (unit-vector) spins via the $S\rightarrow\infty$ limit. This is straightforward in our general-$S$ formalism after spin-operators in Eqns. (1) and (2) are rescaled by $1/S$. The resulting classical coupling and susceptibility are denoted with superscript (c), hence $\chi^{(c)}(\mathbf{k})^{-1}=\underset{S\rightarrow\infty}{\mathrm{lim}}S^{2}\chi(\mathbf{k})^{-1}$ with $X\!\xrightarrow{}\!X^{(c)}/S^{2}$. Note that $b_{1}$ also depends on $S$, hence the right-hand sides of Eqns. (5) to (7) simplify in the limit $S\rightarrow\infty$.
@@
In Fig. 1, we plot the numerical inverse of Eq. (8) [without the unknown contribution $O(X^{4})$] for various $S=1/2$ systems (lines). Note that the lattice only enters via the coordination number $z$. Results are in good agreement with the empirical data points reproduced from the BDMC studies of various $z=4$ and $z=6$ models in Refs. [10, 14, 15].
@@
For all lattices considered here, the corresponding classical temperature is always higher than the quantum temperature. Intuitively, quantum fluctuations heat up the system [10]. However, this is not generally the case with the spin-dimer as a counter-example, see End Matter.
@@
Renormalized MF form of susceptibility.—The QCC is only approximate and breaks down at order $X^{4}$ where first corrections $\epsilon(\mathbf{k})$ to the renormalized MF form appear, see Eq. (7). However, this correction vanishes in the classical limit, $\lim_{S\xrightarrow{}\infty}S^{2}\epsilon(\mathbf{k})=0+{O}((X^{(c)})^{6})$. Therefore, from fourth order on, the momentum dependence of the quantum and classical $\chi(\mathbf{k})$ is inherently different, making it fundamentally impossible to extend the mapping $X(X^{(c)})$ in Eq. (8) for the full $\chi(\mathbf{k})$ to order $X^{4}$. The QCC is therefore always of approximate nature and, contrary to speculations [10], fails for $T\rightarrow 0$.
@@
However, it turns out that corrections $\epsilon(\mathbf{k})$ to the renormalized MF form of $[T\chi(\mathbf{k})]^{-1}$ are relatively small even for small $T/J$ (large $X$). This can be anticipated from the partial cancellation of diagrams as mentioned above, but it can also be inferred empirically by existence of the QCC even for very low $T/J$, e.g. in the pyrochlore case down to $T/J=0.1667$ [14]. Formally, the renormalized MF form of $[T\chi(\mathbf{k})]^{-1}$ is a good approximation to the exact value as long as the minimum of the first two terms $\Delta=f-gz$ (the MF gap) is large compared to $\epsilon(\mathbf{k})$, such that $\|\epsilon\|_{1}/\Delta\ll 1$. Here we use a $L_{1}$ norm in real-space, $\|\epsilon\|_{1}=\sum_{\mathbf{r^{\prime}}}|\epsilon_{\mathbf{r},\mathbf{r^{\prime}}}|$.
@@
In the following we show that for the $d=2,3$ models considered in the context of QCC the susceptibility patterns are very accurately approximated by $\epsilon(\mathbf{k})\approx 0$ well into the cooperative paramagnetic regime $T/J\gtrsim 0.1$. In Fig. 3 we match the $S=1/2$ BDMC data from various models treated in Refs. [10, 14, 15] to the renormalized MF form with the empirically optimized MF temperature $T_{MF}=\frac{fb_{1}}{g}J$. For the $J_{1}\!-\!J_{2}$ models with the renormalized MF form $[T\chi(\mathbf{k})]^{-1}\simeq f+g_{1}\gamma_{1}(\mathbf{k})+g_{2}\gamma_{2}(\mathbf{k})$ (c.f. End Matter), also $J_{2}$ is adjusted, see the labels in Fig. 3. In all instances, the renormalized MF and BDMC data fit very well, even at temperatures below which the analytical mapping based on the truncated version of Eq. (8) would yield reasonable results. This is our second main result. Importantly, this goes beyond the empirical QCC [10] since the exact classical susceptibilities, to which the quantum susceptibilities were matched so far, also feature $O([X^{(c)}]^{6})$ corrections to the renormalized MF form, c.f. the discussion above.
@@
As stated already by Kulagin et al. [10] the QCC clearly fails in the $d=1$ AFM Heisenberg chain. The susceptibility of this model’s classical counterpart is exactly described by the renormalized MF form with $f=\frac{3u^{2}+3}{1-u^{2}}$, $g=\frac{3u}{1-u^{2}}$ and $\epsilon(\mathbf{k})=0$ where $u=\coth(X)-1/X$ [20]: Therefore, QCC can be observed as long as $\|\epsilon\|_{1}/\Delta\ll 1$ for the quantum correlator. We calculate the susceptibility for the $S=1/2$ model with quantum MC [21, 22]. In Fig. 4 we show that already for $T/J=0.33$, the corrections to renormalized MF form are sizable, $\|\epsilon\|_{1}/\Delta\approx 15\%$. This means that the range of $T/J$ where QCC could be observed is significantly smaller in $d=1$ than in the $d=2,3$ models, where no violation was found at the temperatures available to BDMC. This empirical finding can be rationalized in a $1/d$-expansion [17] where corrections to the renormalized MF form at the MF-critical temperature scale as $\epsilon(\mathbf{k})\sim\frac{1}{d^{2}}$.
@@
Figure 4: Parameters of the inverse susceptibility in Eq. (4) for the cyclic $S=1/2$ AFM Heisenberg chain of length $L=128$. Lines are guides to the eye. For better visibility $f,g,\epsilon$ are multiplied by $\chi_{0}\equiv\chi(r=0)$. Breakdown of the renormalized MF form (and of QCC) is rooted in a non-negligible $\|\epsilon\|_{1}/\Delta$.
@@
Figure 5: Static spin structure factor $S(\mathbf{k})=\sum_{\nu,{\nu^{\prime}}}\chi_{\nu,{\nu^{\prime}}}(\mathbf{k})f_{\mathrm{Ni}^{2+}}(\mathbf{k})$ of $\mathrm{K}_{2}\mathrm{Ni}_{2}(\mathrm{SO}_{4})_{3}$ along different planes in reciprocal space, where $f_{\mathrm{Ni}^{2+}}(\mathbf{k})$ is the form factor of the $\mathrm{Ni}^{2+}$ ions. Panels (a)-(c) show classical MC results at $T=0.35J_{4}$ (top left part of each panel), PFFRG results at flow parameter $\Lambda=0.58J_{4}$ (top right), and a fit to the renormalized MF form (rMF, lower half) at $T_{\mathrm{MF}}=0.787J_{4}$. Panels (d)-(f) display line cuts along three principal directions. The classical MC and PFFRG data were taken from Ref. [16].
@@
Finally, we test the renormalized MF form for a model of a complex realistic material $\mathrm{K}_{2}\mathrm{Ni}_{2}(\mathrm{SO}_{4})_{3}$ where QCC was empirically observed by Gonzalez et al. [16]. According to ab-initio calculations [16], this $d\!=\!3$ material realizes a $S\!=\!1$ Heisenberg model on two interconnected trillium lattices with $\{J_{1},J_{2},J_{3},J_{4},J_{5}\}=\{0.066,-0.026,0.144,1,0.479\}$ with dominant $J_{4}=1$. This model was treated with the ground state PFFRG [7] with results extracted at finite Matsubara frequency cutoff parameter $\Lambda=0.58J_{4}$. The latter was argued to act (at least qualitatively) as a finite temperature $T\sim\Lambda$ [23]. Unlike BDMC, the error of PFFRG is uncontrolled. However, empirical confirmation of QCC with $T^{(c)}=0.35J_{4}$ found in Ref. [16] suggests that the susceptibility is well described by a renormalized MF form. In Fig. 5 we match the classical MC and PFFRG susceptibilities [16] along multiple planes and cuts in $\mathbf{k}$-space to a renormalized MF form at $T_{\mathrm{MF}}=0.787J_{4}$ and all $J_{m}$ unchanged. The quality could be further improved by also renormalizing the latter.
@@
Conclusion.—The approximate description of quantum spin systems with classical theories is a recurring theme [24, 25, 26, 27] and cancellations of quantum effects have been observed for various specific spin models on particular lattices [28, 29]. However, the so far empirical QCC for the spin susceptibility on general two- and three dimensional Heisenberg models [10] revisited in this letter stands out by its surprising universality and accuracy even for low temperatures $T$ down to about an order of magnitude below $J$. Using resummed spin diagrammatic perturbation theory, we showed rigorously that the QCC is only approximate and fails at order $(J/T)^{4}$. At this order, the spatial dependence of the exact susceptibility for $S<\infty$ deviates from a simple renormalized MF form.
@@
The success of the renormalized MF form is surprising, since ordinary MF theory is only valid for high temperatures or lattices with large coordination numbers. For future work, it is thus an interesting question if the renormalized MF ansatz could inform the development of novel theoretical (renormalization-)methods or experimental data analysis. The latter could be especially relevant for atom-tweezer array quantum simulators where $\chi(\mathbf{r})$ should be directly accessibly due to single-site control and measurement capabilities [30]. As a first step, it would be useful to extend the analytical (and approximate) QCC mapping of Eq. (8) to higher orders in $(J/T)$ for better estimation of the renormalized MF parameters.
Acknowledgements.
Acknowledgments.—We thank Matías G. Gonzalez, Philip Osterholz, Johannes Reuther, Nepomuk Ritz, Lode Pollet and Manuel Weber for helpful discussions. We acknowledge support from DFG grant no. 524270816. B.Sb. acknowledges support from DFG through the Research Unit FOR 5413/1 (grant no. 465199066). B.Sch. acknowledges funding from the Munich Quantum Valley, supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.